carol ordered 120 air filters for the garage. The number of air filters was 8 less than 2 times the number of gas filters ordered. How many gas filters were ordered?
Answer by gonzo(654)

120 = 2*g -8
solving for g which is number of gas filters gets
2*g = 120 + 8 = 128
g = 128/2
g = 64
120 = 2 * 64 - 8 = 128 - 8 = 120.
64 gas filters is the correct number.